In the lab, it is preferable to measure the volume of a corrosive liquid, instead of weighing it. The reason is you want to avoid ruining an expensive instrument (a balance) if you should spill it. Suppose a corrosive liquid has a density of 1.53 g/mL. What volume would you need to dispense if an experiment calls for 16.1 g of the liquid.


1
Expert's answer
2021-05-26T07:04:17-0400

"Density = \frac{m}{V} \\

1.53 = \frac{16.1}{V} \\

V = \frac{16.1}{1.53}=10.52\;mL"

Answer: 10.52 mL